If 10 men or 20 women or 40 children can do a piece of work in 7 months,
then 5 men 5 women 5 children together can do half of work in?

Answer by richwmiller(17219)

10m 10*7=70 manmonths
10m=20w
20w=40c
c=w/2
w=m/2
5m +5 m/2+ 5 (m/2)/2=(35 m)/4
half of the work is 35 man months
35/((35)/4)=4 months